$SPAD/src/input help.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool help.output
)set message test on
)set message auto off
)clear all
--S 1 of 2
a:= x**2 + 1
--R
--R
--R         2
--R   (1)  x  + 1
--R                                                    Type: Polynomial(Integer)
--E 1
--S 2 of 2
(a - 2)**2
--R
--R
--R         4     2
--R   (2)  x  - 2x  + 1
--R                                                    Type: Polynomial(Integer)
--E 2
)spool
)lisp (bye)

———-

References^{[1] nothing}

4